OpenStudy (anonymous):

The half life of a radioactive substance is 12 years. When will 90% of the material have dissipated?

OpenStudy (schrodingers_cat):

Start with N(t) = No(1/2)^(t/t1/2) t1/2 = half life You want to know when you 90 % of No has dissipated. So, (.1)(No) = No(1/2)^(t/t1/2) Divide by No and take the log of both sides, log(.1) = (t/t1/2)log(1/2) Solve for t, t = (12)log(.1)/(log(1/2) = 39.86 years Hope this helps :)

OpenStudy (anonymous):

@Schrodingers_Cat thank you for aking your time to help me.. but how did you get .1 isn't supposed to be .9 since it is 90%?

OpenStudy (schrodingers_cat):

You want to know when 90% of the material has dissipated. So, only 10% will be left :)

OpenStudy (anonymous):

Oh okay thank you so much @Schrodingers_Cat

OpenStudy (schrodingers_cat):

No Problem :)

OpenStudy (wolf1728):

We can use the formula: elapsed time = half life * log (beginning amount / ending amount) / log 2 elapsed time = 12 * log(100 / 10) / log(2) elapsed time = 12 * 1 / 0.3010299957 elapsed time = 39.8631371386 years You can find half life formulas and a calculator here: http://www.1728.org/halflife.htm